OpenStudy (anonymous):

can i get help proving this identity please tan^2(x)+sin^2(x)+cos^2(x)=sec^2(x)

OpenStudy (anonymous):

tan^2(x) = sec^2(x) - 1 sin^2(x) = 1- cos^2(x) if you substitute those values in for tan^2(x) and sin^2(x) do you see how we can get to the solution?

OpenStudy (anonymous):

the well known identity is sin^2(x)+cos^2(x) = 1. If you divide each term here by cos^2(x) you can derive the identity for tan^2(x) which i wrote earlier

OpenStudy (anonymous):

ok im trying it out

OpenStudy (anonymous):

alright thank u

OpenStudy (anonymous):

Remember that: \(\sin^2(x)+\cos^2(x)=1\). So now we re-write the equation as:\[\tan^2(x)+1=\sec^2(x) \implies \sec^2(x)=sec^2(x) \] \( \therefore L.S=R.S \) @samancha85